If you have an aqueous solution that is 38.3 % Na3PO4 by mass, what is the molality of Na3PO4 in the solution?

Enter your answer in units of molality to three significant figures.
Answer #1

38.3 g of Na3PO4 is present in 100 g of the solution

Mass of the solvent= 100-38.3 = 61.7 g

Molality = moles of the solute/ mass of the solvent in Kg

Moles of Na3PO4= 38.3 g/ 164 g. Mol-1 =0.23353658537 moles

Molality= 0.23353658537 mol/ 0.0617 Kg = 3.79 m